OpenStudy (ashia):

can x^2 + x -12 be factored ?

OpenStudy (kc_kennylau):

(x+4)(x-3)

OpenStudy (ashia):

thanks i got the same answer. was just making sure

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (ashia):

can this one be factored x^2 + 3x + 2 ? i think it cant

OpenStudy (kc_kennylau):

(x+1)(x+2)

OpenStudy (ashia):

how did you get 3x ? lol

OpenStudy (kc_kennylau):

(x+1)(x+2)=x(x+2)+1(x+2)=x^2+2x+x+2=x^2+3x+2

OpenStudy (ashia):

ohh ok i see. thanks so much :)

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (ashia):

Simplify/Reduce the Algebraic Fraction x^2 + 2x + 1 /3x^5 -x^4

OpenStudy (ashia):

i got 3x -1

OpenStudy (ashia):

oops i typed it wrong its x^2 + 2x + 1 /x+1

OpenStudy (kc_kennylau):

nope it's not correct, factorize the numerator first :)

OpenStudy (kc_kennylau):

\[\frac{x^2+2x+1}{x+1}=\frac{(x+1)(x+1)}{x+1}=x+1\]

OpenStudy (kc_kennylau):

sorry gtg cya WolframAlpha can help you check your answers bye have a merry Christmas

OpenStudy (ashia):

ok no problem :) thank you for your help. merry Christmas and a happy new year :)